# Anatomical Foundation Models for Brain MRIs

### v3

@@ 4.7 Ablation study
We perform an ablation on different neural architectures, considering larger models such as 3D ResNet-50. We report the results of this ablation on the Schizconnect dataset here in Tab. 6. From the obtained results, we do not find significant advantages in using a 3D ResNet-50 over the simpler and smaller version 3D ResNet-18. For this reason, we chose to employ this simpler architecture, as it is practically more convenient to train and deploy in real settings. It is worth mentioning that similar findings are reported in related literature [13], showing that deeper architectures such as even a 3D DenseNet-121 do not bring significant improvements over a 3D ResNet-18, at least on VBM-processed images. We did not explore different and larger architectures such as ViT [11], as we were limited by computational constraints. In fact, a 3D ViT implementation requires significantly more memory and computational power, which makes it less practical and favorable with respect to ResNet-18, also from a carbon footprint point of view.
@@
ResNet-18 y-Aware [12] 66.20$\pm$4.50 71.04$\pm$2.31 70.29$\pm$14.73 63.95$\pm$19.81
ExpW [1] 69.53$\pm$4.43 67.65$\pm$8.27 63.26$\pm$18.06 74.34$\pm$18.95
AnatCL-G3 72.55$\pm$5.16 71.03$\pm$8.53 73.65$\pm$7.29 63.43$\pm$15.94
AnatCL-L3 66.38$\pm$5.96 72.07$\pm$8.42 68.37$\pm$8.60 56.60$\pm$11.48
@@
AnatCL-G3 68.07$\pm$4.16 72.16$\pm$5.53 57.92$\pm$14.50 78.71$\pm$18.67
AnatCL-L3 65.20$\pm$3.19 62.99$\pm$8.39 69.40$\pm$9.98 67.61$\pm$15.81
Table 6: Model size ablation study. Larger architectures do not result in consistent improvements over simpler models, as also shown in [13].
@@
5 Conclusions
We propose AnatCL, a foundation model based on brain anatomy and trained with a weakly contrastive learning approach. With thorough validation on 10 different downstream tasks, we show that incorporating anatomy information during training can results in more accurate predictions of different neurological and psychiatry conditions, and also partially clinical assessment scores and phenotypes. We release the weights of the trained model for public use at omittedlink, empowering researchers and practitioners worldwide to leverage AnatCL in many different applications.
Supplementary Material
6 Limitations, impact, and ethics
7 Results
@@ 7.1 Assessment scores and phenotypes
Table 8: Assessment scores / phenotypes prediction from brain MRIs.

### v4

@@ 7.2 Ablation study
We perform an ablation on different neural architectures, considering larger models such as 3D ResNet-50. We report the results of this ablation on the Schizconnect dataset here in Tab. 6. From the obtained results, we do not find significant advantages in using a 3D ResNet-50 over the simpler and smaller version 3D ResNet-18. For this reason, we chose to employ this simpler architecture, as it is practically more convenient to train and deploy in real settings. It is worth mentioning that similar findings are reported in related literature (Dufumier et al., 2022), showing that deeper architectures such as even a 3D DenseNet-121 do not bring significant improvements over a 3D ResNet-18, at least on VBM-processed images. We did not explore different and larger architectures such as ViT (Dosovitskiy, 2020), as we were limited by computational constraints. In fact, a 3D ViT implementation requires significantly more memory and computational power, which makes it less practical and favorable with respect to ResNet-18, also from a carbon footprint point of view.
Table 6: Model size ablation study. Larger architectures do not result in consistent improvements over simpler models, as also shown in Dufumier et al. (2022).
@@
ResNet-18 y-Aware 66.20$\pm$4.50 71.04$\pm$2.31 70.29$\pm$14.73 63.95$\pm$19.81
ExpW 69.53$\pm$4.43 67.65$\pm$8.27 63.26$\pm$18.06 74.34$\pm$18.95
AnatCL-L3 66.03$\pm$4.06 70.03$\pm$7.16 72.47$\pm$12.27 63.18$\pm$21.50
AnatCL-G3 64.31$\pm$4.43 73.04$\pm$4.42 62.42$\pm$14.95 84.19$\pm$7.10
@@
AnatCL-G3 (Desikan) 68.07$\pm$4.16 72.16$\pm$5.53 57.92$\pm$14.50 78.71$\pm$18.67
AnatCL-L3 (Desikan) 65.20$\pm$3.19 62.99$\pm$8.39 69.40$\pm$9.98 67.61$\pm$15.81
@@ 7.3 Assessment scores and phenotypes
Table 8: Assessment scores / phenotypes prediction from brain MRIs.